# GEIC: Universal and Multilingual Named Entity Recognition with Large Language Models

### v2

@@ 3.2.1 Extraction
In the extraction sub-task, we utilize a generation-based extraction method, where special tokens "##" are used to surround any entities identified in the sentence, regardless of the number of entities or their types For example:
Query: Kobe was in NBA
Response: ##Kobe## was in ##NBA##
@@
This method, compared to conventional sequential extraction, avoids requiring LLMs to perform text alignment, thus reducing task complexity. In contrast to similar methods Wang et al. (2023); Hu et al. (2024b), CascadeNER’s query contains only the sentence, without any task descriptions, demonstrations, or category information. The response exclusively uses "##" as the identifiers, and all entities are extracted without specifying categories. CascadeNER achieves low-cost NER by using simple prompts and better generalization by treating all entities uniformly. Moreover, extracting in the sentence-level leverages LLMs’ self-attention mechanisms for global understanding, improving accuracy in low-context scenarios. A detailed comparison with existing methods and further advantages of our method are shown in Appendix D.
@@
where $E_{i}$ is the set of extracted entities from the $i$-th extraction, $n$ is the number of extraction rounds, $E_{\text{overlap}}$ is the set of overlapping or nested entities, and $\text{length}(e)$ is the length of entity $e$. The effects of the number of extraction repetitions and other details are provided in Appendix C.1.
3.2.2 In-context Classification
@@
Figure 3: Examples of classification.
For multi-granularity data, we apply hierarchical classification, ensuring the accuracy of classification for complex categories. After obtaining the coarse-grained result, CascadeNER use the result to index the corresponding sub-categories and classify again, continuing this process until no further classification is possible:
@@
In zero-shot scenarios, we use a slightly different prompt. Due to significant differences in entity classifications across datasets, some named entities in one dataset may be overlooked in others. We append the query with If none of them applied, return unknown to handle situations where the extracted entity cannot be classified into the provided categories, enhancing CascadeNER’s generalization across datasets.
3.2.3 Few-shot Data Sampling
In existing datasets, only CrossNER Liu et al. (2021), designed for low-resource scenarios, and FewNERD Ding et al. (2021), designed for few-shot scenarios, meet our requirements for evaluating CascadeNER in few-shot scenarios. However, relying solely on them is insufficient for comprehensively evaluating CascadeNER, particularly its multilingual NER performance.
@@ 3.2.3 Few-shot Data Sampling
To address this, we develop a sampling algorithm to construct datasets for few-shot evaluation. Considering that basic random sampling cannot ensure a balanced category distribution, we employ a stratified sampling algorithm, which divides the dataset into strata based on the labels. Each stratum corresponds to a distinct entity type, and we ensure an relatively equal number of samples per category by drawing from these strata, thereby maintaining balance across categories in the results. The size for each stratum is calculated with the formula:
@@
4 AnythingNER Dataset
Due to the tremendous parameters and the self-attention mechanism, LLMs demonstrate uniquely high performance in low-resource scenarios for NER tasks. However, there are no existing NER datasets optimized for the characteristics of LLMs, limiting the comprehensive training and evaluation of LLMs in these tasks and making it difficult to effectively evaluate GEIC. Moreover, there is also a lack of multilingual and fine-grained datasets.
Figure 4: The 8 coarse-grained categories, 31 medium-grained categories of AnythingNER. More detailed categories are provided in Appendix G
We thus constructed AnythingNER, the first NER dataset specifically designed for LLM-based NER, particularly GEIC tasks. AnythingNER spans 8 major languages: English, Chinese, Spanish, French, German, Japanese, Korean, and Russian. In terms of classification, it is the first NER dataset to feature three-level granularity categorization, encompassing 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, as illustrated in Figure 4.
Compared to existing datasets, AnythingNER offers several key advantages. It adopts a three-level granularity categorization with 155 categories, ensures comprehensive coverage across diverse domains, including rare entities and emerging fields like technology products, which are not adequately considered in earlier datasets. It is specifically formatted to be compatible with GEIC tasks and supports 8 major languages to leverage LLMs’ multilingual capabilities. Additionally, the dataset ensures balanced category distribution for few-shot scenarios and applies a novel dynamic categorization for zero-shot scenarios, making it highly suitable for current research focuses.
Figure 5: Example of GEIC format.
AnythingNER provides data in three formats: the standard BIO format, a novel format for GEIC tasks termed GEIC format, and the dynamic categorized GEIC format. The BIO format follows the same structure as classic datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). The structure of the GEIC format is shown in Figure 5. This format is easy to index and includes language information, facilitating data processing, especially in multilingual scenarios. The standard GEIC version is designed for few-shot scenarios, while the dynamic categorized version is for zero-shot scenarios. The core of our dynamic categorization is dynamically adjusting and simplifying the entity type list provided to the model for improving the model’s generalization and reducing the risk of overfitting. Detailed implementation and advantages are provided in Appendix A.
All the corpora in AnythingNER are sourced from Wikipedia contributors (2024) and manually annotated by our team. A statement regarding the annotators and ethical considerations can be found in Appendix F. We also provide some existing open-source datasets converted into GEIC format. All versions of AnythingNER are divided into train, dev, and test sets. Further details, including data volumes for different languages, are provided in Appendix B.
5 Experiment
5.1 Experiment Setting
Model English Spanish French Russian German Chinese Japanese Korean
XLM-RoBERTa 41.9 33.5 29.1 23.4 32.9 29.2 27.2 28.6
GEMNET 45.1 37.3 38.4 33.4 37.2 40.7 33.2 35.4
LS-unLLaMA 52.3 47.5 42.1 30.7 35.6 38.0 39.5 41.8
GPT-NER 46.3 40.4 38.8 34.8 39.7 35.1 36.3 36.1
CascadeNER-1.5B 58.2 52.3 51.5 49.8 47.7 54.5 52.7 51.3
CascadeNER-combine 60.2 54.2 53.4 50.5 49.1 57.4 53.2 52.5
CascadeNER-7B 60.3 55.1 54.0 51.7 50.3 58.8 54.7 52.9
Table 1: F1 score of different models across languages in the standard BIO and GEIC AnythingNER.
@@ 5.1 Experiment Setting
Implementation
In our experiments, we use Qwen2-1.5B and Qwen2-7B Yang et al. (2024) as the base models for CascadeNER, as they represent the current best-performing SLMs. We also evaluate Gemma, another prominent SLM, as detailed in Appendix C.2. Based on the dynamic categorized version of AnythingNER, we utilize the SWIFT framework Team (2024) to perform supervised fine-tuning (SFT) on the instruct versions of these models. Each model is fine-tuned separately on the corresponding dataset to obtain an extractor and a classifier. In addition to model sets where the extractor and classifier are of the same size, we also experiment with a configuration where the extractor uses the 1.5B model and the classifier uses the 7B model, aiming to balance accuracy and efficiency since the classification task is more complex. Due to page constraints, in experiments outside of AnythingNER, we consistently use the Qwen2-1.5B model in CascadeNER, emphasizing the framework’s low computational resource requirements and its suitability for SLMs. We evaluate CascadeNER’s performance in both few-shot and zero-shot scenarios, comparing it with supervised SOTAs and LLM-based baselines. The repetition count $i$ for result fusion is set to 3.
Dataset
We conduct experiments on the AnythingNER and existing datasets including CoNLL2003 Tjong Kim Sang and De Meulder (2003), CrossNER Liu et al. (2021), FewNERD Ding et al. (2021), PAN-X Pan et al. (2017), and MultiCoNER Malmasi et al. (2022). Since we decide to use and share these datasets in the GEIC format, we only choose open-sourced datasets to avoid copyright issues. For evaluation metrics, we primarily use F1 score, as it is widely recognized as the most robust and effective metric for NER tasks Li and Sun (2020). Potential data contamination are discussed in Appendix E.
Baselines
For supervised methods, we adopt ACE+document-context by Wang et al. (2020) (SOTA of CoNLL2003) and BERT-MRC+DSC by Li et al. (2019b) (SOTA of Ontonotes 5.0 Pradhan et al. (2013)) for English datasets, while XLM-RoBERTa Conneau et al. (2020) and GEMNET by Meng et al. (2021); Fetahu et al. (2022) (SOTA of MultiCoNER) for multilingual datasets. For LLM-based methods, we adopt GPT-NER with GPT-3.5 by Wang et al. (2023), LS-unLLaMA with LLaMA2-7B by Li et al. (2023), and PromptNER with GPT-4 by Ashok and Lipton (2023) (the current SOTA of CrossNER and FewNERD). These methods achieve the best performance when paired with the respective models according to their original papers. GPT-NER and PromptNER are few-shot learning methods, while LS-unLLaMA is a supervised method. All few-shot samples are sampled using the algorithm in Section 3.2.3. Due to the expensive price of GPT, we also applied this algorithm to limit the test set size to 300 sentences for all models. For few-shot demonstrations, PromptNER is provided with 2-shot for the best performance according to its paper, while GPT-NER and CascadeNER provided with 50-shot. Notably, when provided with more demonstrations, the performance of GPT-NER and CascadeNER improves further. We provide CascadeNER’s best performance results under the 100-shot and dual 7B model configuration in Appendix H.
@@
5.2 Results
AnythingNER
The results are shown in Table 1. Notably, since the base models of CascadeNER are fine-tuned with AnythingNER, we only provide zero-shot results. The results indicate that, due to its unprecedentedly detailed categorization and multilingual coverage, AnythingNER is a extremely challenging flat NER dataset, placing higher demands on methods’ generalization. Additionally, its low-resource characteristics make AnythingNER more suitable for evaluating few-shot and zero-shot methods, particularly LLM-based methods. CascadeNER achieve a significant leading advantage on this dataset, demonstrating its strong generalization and multilingual proficiency. PromptNER experience serious hallucinations when faced with the complete set of 155 categories, so we fail to provide results of it.
@@ 5.2 Results
MultiCoNER and PAN-X are two widely used multilingual datasets. MultiCoNER covers 6 entity types and 11 languages, while PAN-X includes 3 entity types and 282 languages. We use 6 and 8 overlapping languages from MultiCoNER and PAN-X with AnythingNER to evaluate CascadeNER’s multilingual capabilities. The results are shown in Table 3. It is important to note that, for the purpose of controlling variables, all methods requiring training, including CascadeNER’s base model pre-training, are trained using multilingual joint training.
Discussion and Limitation

### v4

@@ 4.2 Extraction
In the extraction sub-task, we utilize a generation-based extraction method, where special tokens "##" are used to surround any entities identified in the sentence, regardless of the number of entities or their types. For example:
Q: Apple proposes new Macbook
A: ##Apple## proposes new ##Macbook##
@@
This method, compared to conventional sequential extraction, avoids requiring LLMs to perform text alignment, thus reducing task complexity. Comparing similar methods Wang et al. (2023); Hu et al. (2024b), CascadeNER’s query contains only the sentence, without any task descriptions, demonstrations, or category information. The response exclusively uses "##" as the identifiers, and all entities are extracted without specifying categories. CascadeNER achieves low-cost NER by using simple prompts and better generalization by treating all entities uniformly. A detailed comparison with existing methods and further advantages of our method are shown in Appendix F.
Result Fusion
@@
where $E_{i}$ is the set of extracted entities from the $i$-th extraction, $n$ is the number of extraction rounds, $E_{\text{overlap}}$ is the set of overlapping or nested entities, and $\text{length}(e)$ is the length of entity $e$. The effects of the number of extraction repetitions and other details are provided in Appendix E.1.
4.3 Classification
Prompt Design
@@ 4.3 Classification
Figure 5: Example prompts of classification.
In zero-shot scenarios, we use a slightly different prompt. Due to differences in entity categorization across datasets, some entities in one dataset may be overlooked in others. We append the query with If none of them applied, return unknown to handle situations where the extracted entity cannot be classified into the provided categories, enhancing CascadeNER’s generalization.
Multi-granularity
For multi-granularity data, we apply a progressive strategy. After obtaining the coarse-grained result, CascadeNER use the result to index the corresponding sub-categories and classify again, continuing this process until no further classification is possible:
@@
5 Experiment
In this section, we first present the categorization metric changes of DynamicNER before and after dynamic categorization, followed by a comparative analysis of existing methods and CascadeNER’s performance on different versions of DynamicNER. We also conduct experiments of CascadeNER and baselines on existing datasets, along with ablation studies. In evaluations across existing datasets, CascadeNER, with base models fine-tuned using the dynamic version of DynamicNER, demonstrates consistent excellence in all datasets and achieves new SOTA performance in both FewNERD and CrossNER datasets. This confirms that DynamicNER not only provides exceptional effectiveness for evaluating LLM-based NER methods but also offers substantial value in training. Due to space limitations, these experiments are provided in Appendix D and E.
Figure 6: Quantitative categorization metric results for 3 versions DynamicNER in English and Chinese. Generally, higher cohesion and normalized entropy, or lower Gini coefficient and variation coefficient, indicate better quality.
5.1 Categorization Quality Evaluation
To demonstrate that our dynamic categorization improves dataset generalization while maintaining dataset quality, we conduct comparative experiments across three versions of DynamicNER: the Base Version, a version with random parameters for dynamic categorization, and the optimized Dynamic Version. We still employ the 4 metrics for evaluating dataset quality, whose detailed definition are provided in Appendix A. For reliability of the random version’s results, we conduct five independent tests and use the average results. Due to space limitations, we only present results for English and Chinese in Figure 6 here. Other quantitative results are provided in Appendix K.
Experimental results demonstrate that our dynamic categorization significantly increases data diversity, as shown in Table 2, while maintaining or improving dataset quality compared to the base version. The quality of the dynamic version also considerably surpasses the random version. These results comprehensively validate the reliability and effectiveness of our method.
Language en es fr ru de zh ja kr
# Lists 725 455 501 377 465 786 553 478
Table 2: The numbers of entity type lists of each language after dynamic categorization. In some scenarios, this can be equivalent to having 700+ distinct datasets.
5.2 DynamicNER Experiment
Baseline Selection
In our experiments for DynamicNER, we evaluate four NER methods: XLM-RoBERTa Conneau et al. (2020), GPT-NER Wang et al. (2023), PromptNER Ashok and Lipton (2023), and our CascadeNER. XLM-RoBERTa is a famous BERT-based multilingual model widely used as a baseline in multilingual NER research Malmasi et al. (2022); Fetahu et al. (2022), thus being selected as our baseline representing supervised methods. GPT-NER and PromptNER are two major general LLM-based NER methods that achieve performance significantly superior to supervised methods in low-resource scenarios through sophisticated prompt design and powerful GPT models, as discussed in Section 2 and Appendix F.
Model Selection
Given the lack of existing lightweight LLM-based NER methods and controlled variable principles, we evaluate two LLM-based methods and CascadeNER using three LLMs: Qwen2.5-1.5B, Qwen2.5-7B, and GPT-4o. Qwen series is the largest open-source model family. Its lightweight LLMs perform exceptionally across benchmarks and gaining widespread recognition. According to HuggingFace (2024), Qwen2.5-1.5B is the most downloaded open-source model in 2024. Therefore, we select Qwen2.5-1.5B and 7B to represent the current best-performing lightweight LLMs. GPT-4o is the most widely-used current general commercial LLM, and its previous versions are employed in GPT-NER and PromptNER, making it our representative choice for commercial LLMs. In CascadeNER, for statistics convenience, the extractor and the classifier use the same base model.
Model Dynamic-Supervised Dynamic-Fewshot
en es fr ru de zh ja ko en es fr ru de zh ja ko
G-1.5B 47.6 39.7 38.0 37.6 37.3 41.2 35.7 36.1 36.9 32.2 31.9 30.5 30.3 35.8 31.9 32.6
G-7B 52.3 46.4 44.8 44.8 45.7 48.1 42.3 42.1 42.7 37.3 38.2 36.8 36.5 41.1 37.3 38.6
G-GPT 60.6 57.3 56.5 55.6 55.9 58.4 54.9 53.8 49.2 46.9 47.5 47.2 47.0 48.9 47.7 48.3
P-1.5B 23.2 20.8 18.5 16.3 17.5 22.7 18.0 17.3 20.5 17.9 16.2 15.9 16.1 19.9 16.0 15.9
P-7B 44.3 35.8 33.2 32.5 31.9 40.4 37.4 35.6 39.8 33.2 32.1 31.8 31.5 37.8 35.6 34.5
P-GPT 53.0 50.5 51.2 47.9 50.2 52.3 48.7 48.5 49.4 48.5 47.1 46.6 46.0 47.4 44.1 44.0
C-1.5B 62.8 55.7 52.8 51.1 48.8 58.9 54.1 52.7 49.7 44.1 44.0 43.4 42.9 48.5 43.1 43.8
C-7B 68.2 61.5 55.3 52.9 51.4 64.5 58.8 55.3 55.7 49.9 49.7 46.5 46.1 52.9 50.2 50.0
C-GPT 73.1 67.1 67.8 66.9 67.6 68.3 67.4 67.9 61.3 57.4 56.9 56.2 56.0 59.7 56.8 56.4
Table 3: The results of supervised learning with dynamic version and few-shot learning with dynamic version. G means GPT-NER, P means PromptNER, and C means CascadeNER. The results indicate that, due to its unprecedentedly detailed categorization and multilingual coverage, DynamicNER is a extremely challenging flat NER dataset, placing higher demands on methods’ generalization capability.
Implementation
For the supervised method, XLM-RoBERTa is trained and only evaluated with the base version of DynamicNER. As its fixed classification output layer corresponds to a predefined set of entity types and any modification to the entity type list necessitates full model retraining, it can not be evaluated with the dynamic version. For LLM-based methods, we conduct experiments under three scenarios: supervised learning with base version, supervised learning with dynamic version, and few-shot learning with dynamic version. Training data for GPT-NER and CascadeNER is obtained through format conversion. For PromptNER, as its prompt involves complex designs such as CoT, we utilize LLM-generated prompts by GPT-4o using prompts from its paper as few-shot demonstrations and manually verified the prompts. The repetition count $i$ of CascadeNER for result fusion is set to 3. Potential data contamination are discussed in Appendix H.
Model Base-Supervised
en es fr ru de zh ja ko
BERT 41.9 33.5 29.1 23.4 32.9 29.2 27.2 28.6
G-1.5B 50.2 43.5 40.4 39.8 39.3 44.1 38.9 38.7
G-7B 55.1 48.2 47.2 44.0 48.1 50.9 44.8 44.5
G-GPT 62.4 58.3 57.9 56.8 56.9 60.4 57.3 55.9
P-1.5B 21.6 18.6 17.1 14.9 15.8 20.7 16.4 15.9
P-7B 41.1 32.9 31.0 30.7 30.3 47.4 35.6 29.6
P-GPT 49.7 47.7 48.2 45.9 46.6 48.6 45.7 45.4
C-1.5B 67.6 59.9 57.9 55.7 53.5 64.0 58.5 55.1
C-7B 73.8 65.5 60.3 59.6 61.4 69.8 65.3 62.7
C-GPT 77.1 71.7 69.9 70.3 70.8 74.3 72.4 70.9
Table 4: The results of supervised learning with base version. RoBERT represents XLM-RoBERTa. The supervised method XLM-RoBERTa performs terribly.
Results and Discussion
The results are shown in Table 3 and 4. CascadeNER achieve a significant advantage on DynamicNER, demonstrating its strong generalization and multilingual proficiency. The supervised method XLM-RoBERTa performs terribly, as DynamicNER’s low-resource characteristics make it more suitable for evaluating LLM-based methods. For LLM-based methods, the 3 methods show significant performance variations across different datasets and models. When using GPT-4o and transitioning from supervised to few-shot, PromptNER exhibits notably smaller performance degradation, partially reflecting the generalization advantages of reasoning-focused approaches. However, when migrating to lightweight LLMs, these methods show significantly larger performance drops compared to the other two methods. GPT-NER and CascadeNER demonstrate generally similar performance patterns, but GPT-NER shows more pronounced degradation when migrated to lightweight LLMs, while CascadeNER achieves a greater performance advantage on the dynamic version compared to the base version, validating the effectiveness of the two-stage strategy in both complex classification and reducing computional resource requirement.
6 Conclusion
This paper introduces DynamicNER, a multi-lingual and multi-granular NER dataset optimized for LLM-based NER, including a human-annotated base version and a dynamic-categorized version. We develop the first dynamic categorization method in NER datasets for DynamicNER, enhancing its generalization while keeping data quality. We also propose CascadeNER, a powerful NER framework which is exceptionally suitable for lightweight LLMs and local deployment, outperforming current LLM-based methods. Moreover, we conduct comprehensive experiments and analyses on DynamicNER and discuss the advantage and future direction of LLM-based NER. More experiments and discussions are provided in Appendix D and F. We hope that DynamicNER and CascadeNER will facilitate future research in LLM-based NER, revitalizing this classical NLP task.
7 Limitation
There are still some challenges for our research. Although CascadeNER is designed to be able to accommodate nested and discontinuous NER, we only conduct evaluation on CascadeNER about flat NER tasks. This limitation arises from the fact that the models in CascadeNER are pre-trained on the dynamic version of DynamicNER, and DynamicNER is a flat NER dataset. Our resources are insufficient to collect enough open-source data for this purpose, which lead to DynamicNER containing only flat NER labels, and thus constraining CascadeNER to flat NER. Furthermore, Due to resource constraints and our failure to find annotators proficient in other languages for manual annotation, DynamicNER currently supports only 8 languages, which somewhat restricts its applicability.
Appendix A Categorization Metric Definition
Cohesion
Category cohesion score (cohesion) measures categorical semantic consistency by calculating the average semantic similarity between all entities within the same category. We employ the BERT-base Devlin et al. (2018) model to extract semantic representations of entities, obtain embeddings of each entity, then computing cosine similarity between embeddings to derive cohesion. This metric ranges from [-1,1], where 1 indicates complete similarity and -1 indicates complete opposition. Typically, we perform category merging when cohesion exceeds 0.9. The formula is shown below:
$Cohesion=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j=i+1}^{n}\cos(\mathbf{v}_{i},\mathbf{v}_{j})$ (3)
where $n$ is the number of entities in this category, $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ are the vector representations of the $i$-th and $j$-th entities encoded by BERT-base, and $\cos(\mathbf{v}_{i},\mathbf{v}_{j})$ represents the cosine similarity between two vectors.
The detailed cosine similarity formula is shown below.
$\cos(\mathbf{v}_{i},\mathbf{v}_{j})=\frac{\mathbf{v}_{i}\cdot\mathbf{v}_{j}}{\|\mathbf{v}_{i}\|\|\mathbf{v}_{j}\|}$ (4)
Normalized Entropy
Normalized entropy measures the overall balance of category distribution. This metric is used for the influence of category quantity by calculating the information entropy of category frequency distribution and normalizing it to a score within the range [0,1]. A score of 1 indicates perfect balance, where all categories have equal sample sizes, while 0 indicates complete imbalance, where all samples are concentrated in a single category. When normalized entropy falls below 0.8, it indicates significant distributional imbalance and needs to be adjusted. The formula is shown below:
$H=-\frac{\sum_{i=1}^{n}p_{i}\log_{2}(p_{i})}{\log_{2}(n)}$ (5)
where $n$ is the total number of categories, $p_{i}$ is the proportion of samples in the $i$-th category calculated as the number of samples in category $i$ divided by the total number of samples across all categories.
Gini Coefficient
The Gini Coefficient Gini (1921) measures the degree of inequality in category distribution. Compared to normalized entropy, the Gini coefficient demonstrates higher sensitivity to distributional inequalities, performing better at identifying extreme imbalances where minority categories contain large sample proportions. For instance, when sample distributions exhibit extreme imbalances like [0.8, 0.1, 0.05, 0.05], the Gini coefficient provides stronger warning signals, while normalized entropy is more suitable for monitoring progressive imbalances such as [0.4, 0.3, 0.2, 0.1]. This metric also ranges from [0,1], where 0 indicates perfect balance and 1 indicates complete imbalance. A Gini coefficient exceeding 0.4 signals significant categorical inequality and requires distribution improvement. By using Gini coefficient and normalized entropy together, we achieve both sensitive detection of extreme imbalances and effective monitoring of overall distribution trends. The formula is shown below:
$G=\frac{n+1-2\sum_{i=1}^{n}(n-i+1)p_{i}}{n}$ (6)
where $n$ is the total number of categories, $p_{i}$ is the proportion of samples in the $i$-th category after sorting proportions in ascending order $(p_{1}\leq p_{2}\leq...\leq p_{n})$.
Variation Coefficient
The Coefficient of Variation measures data dispersion by calculating the ratio of standard deviation to mean of category sample sizes. Its advantage is its scale independence, enabling comparisons across different scenarios. The coefficient ranges from 0 to positive infinity, where 0 indicates perfect balance and larger values indicate greater distributional imbalance. When the coefficient exceeds 0.5, it indicates significant fluctuation in sample sizes between categories, necessitating balance adjustments. The formula is shown below:
$CV=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}}{\bar{x}}=\frac{\sigma}{\mu}$ (7)
where $n$ is the total number of categories, $x_{i}$ is the number of samples in the $i$-th category, $\bar{x}$ is the mean number of samples across categories, $\sigma$ is the standard deviation of sample numbers, and $\mu$ is the mean.
Appendix B Categorization Metric Selection
Mixing Types of Different Granularities
In this round of re-categorization, we use cohesion, normalized entropy, and Gini coefficient as metrics for optimization. Cohesion is employed to assess relationships between entity types, where close categorical relationships reduce the need for mixing to avoid creating unreasonable combinations. Meanwhile, normalized entropy and Gini coefficient are utilized to comprehensively measure distribution uniformity, where uneven distributions guide the system to perform additional merging for balance or category redistribution.
Replace with Synonyms
In this round of re-categorization, we use Gini coefficient and variation coefficient as metrics for optimization. We employ the variation coefficient to measure data dispersion, increasing synonym substitutions for increasing data convergence when dispersion is high. The Gini coefficient is used to guide system to reduce operations to prevent exacerbating imbalances when distributions are uneven. Cohesion is not used as synonym substitution does not alter hierarchical relationships between categories. Entropy is also given up because synonym substitution primarily focuses on linguistic variation rather than distributional changes.
Remove Irrelevant Types
In this round of re-categorization, we use cohesion and normalized entropy as metrics for optimization. We employ normalized entropy as a reference for controlling removal probability, ensuring that deletion operations do not result in overly concentrated distributions. Additionally, the system adjusts removal probability when cohesion is low, regulating relationships between categories. The variation coefficient is not used as this stage primarily focuses on option quantity rather than distribution characteristics, while the Gini coefficient is omitted since distribution balance has been addressed in the previous two stages, thus temporarily foregoing the Gini coefficient to prevent interference with other metrics.
Merge Types into Miscellaneous
In this round of re-categorization, we use all four metrics for optimization. As the final optimization stage, it requires consideration across all dimensions. We use all metrics for final fine-tuning to ensure overall data quality and avoid biases that might arise from single metric optimization.
Appendix C Detail of DynamicNER
The specific data volumes for each language are shown in Table 5. It is important to note that for languages except English and Chinese, we partially use manually translated English corpora. This is necessary to balance category distribution, as some languages lack sufficient corpora in specific domains. We also provide conversion scripts that allow DynamicNER to be transformed into train, dev, and test sets with non-overlapping subsets based on coarse categories, making it easier to use for traditional few-shot learning methods.
Additionally, two points about the Table 5 require clarification. First, as DynamicNER’s design emphasizes the evaluation of generalization and low-resource learning capabilities, we set the test set capacity to the biggest one, rather than the train set. Second, for Chinese, Japanese, and Korean, due to linguistic characteristics where each character is treated as a token, the token count appears significantly higher, though the actual corpus volume is comparable to other languages.
Language # Sentences # Tokens # Entities # Train # Dev # Test
English 1500 36.7k 4664 300 300 900
Chinese 1500 98.1k 5198 300 300 900
Spanish 1000 22.8k 2454 197 201 602
French 1000 24.1k 2763 200 200 600
German 1000 21.7k 2800 200 197 603
Japanese 1000 81.7k 3032 201 199 600
Korean 1000 66.4k 2401 202 200 598
Russian 1000 18.5k 2092 201 198 601
Table 5: Statistics of DynamicNER across languages. We roughly follow a 1:1:3 ratio to divide the train, dev, and test sets, with slight adjustments based on the proportional distribution of entities within the corpus.
Appendix D More Experiment about CascadeNER
D.1 CascadeNER Setting
In the experiments of this section, CascadeNER always employs two Qwen2.5-7B base models, which are fine-tuned separately based on the corresponding part of the dynamic version of DynamicNER to obtain an extractor and a classifier. Potential data contamination about the fine-tuning is discussed in Appendix H. We evaluate CascadeNER’s performance in both few-shot and zero-shot scenarios, comparing it with supervised SOTAs and LLM-based baselines. For few-shot scenarios, the number of few-shot demonstrations is set to $3$, the same as the experiments on DynamicNER.
D.2 Baselines
For supervised methods, we adopt ACE+document-context by Wang et al. (2020) (SOTA of CoNLL2003) and BERT-MRC+DSC by Li et al. (2019b) (SOTA of Ontonotes 5.0 Pradhan et al. (2013)) for English datasets, while XLM-RoBERTa Conneau et al. (2020) and GEMNET by Meng et al. (2021); Fetahu et al. (2022) (SOTA of MultiCoNER) for multilingual datasets. For LLM-based methods, we adopt GPT-NER and PromptNER with GPT-4o.
D.3 Dataset
Few-shot Data Sampling
@@ D.3 Dataset
In existing datasets, only CrossNER Liu et al. (2021), designed for low-resource scenarios, and FewNERD Ding et al. (2021), designed for few-shot scenarios, meet our requirements for evaluating CascadeNER in few-shot scenarios. However, relying solely on them is insufficient for comprehensively evaluating CascadeNER, particularly its multilingual NER performance. To address this, we develop a sampling algorithm to construct datasets for few-shot evaluation. Considering that basic random sampling cannot ensure a balanced category distribution, we employ a stratified sampling algorithm, which divides the dataset into strata based on the labels. Each stratum corresponds to a distinct entity type, and we ensure an relatively equal number of samples per category by drawing from these strata, thereby maintaining balance across categories in the results. The size for each stratum is calculated with the formula:
@@
Dataset Selection
We conduct supplementary experiments on existing datasets including CoNLL2003 Tjong Kim Sang and De Meulder (2003), CrossNER Liu et al. (2021), FewNERD Ding et al. (2021), PAN-X Pan et al. (2017), and MultiCoNER Malmasi et al. (2022). Since we decide to use and share the formatted versions of these datasets in our repository to facilitate the test and use of CascadeNER, we only choose open-sourced datasets to avoid copyright issues. For evaluation metrics, we primarily use F1 score, as it is widely recognized as the most robust and effective metric for NER tasks Li and Sun (2020). We detail below the reasons for selecting these datasets and their usage.
@@
MultiCoNER and PAN-X are two widely used multilingual datasets. MultiCoNER covers 6 entity types and 11 languages, while PAN-X includes 3 entity types and 282 languages. We use 6 and 8 overlapping languages from MultiCoNER and PAN-X with DynamicNER to evaluate CascadeNER’s multilingual capabilities. It is important to note that, for the purpose of controlling variables, all methods requiring training are trained using multilingual joint training.
D.4 Experimental Results